# Jet-Driven Formation of Bipolar Rings in Planetary Nebulae: Numerical Simulations Inspired by NGC 1514

### v1

@@ 3.2.4 Effects of Radiative Cooling
Figure 10 of the column density maps for $\gamma=1.05$ presents at late times a bipolar ring structure, but not as prominent as in the adiabatic $\gamma=5/3$ simulations that we present in Figures 9 and 11. To better assess the impact of radiative cooling on the ring formation process, we performed an additional simulation including optically thin radiative cooling using the tabulated cooling function of Sutherland & Dopita (1993); we kept all other initial conditions and jet parameters as in the adiabatic simulation with the thick dense shell. In Figure 13 we present the column density of this simulation at $t=220{~\rm yr}$ and an inclination angle of $i=50^{\circ}$.
Figure 13: A column-density map of a simulation with a radiative cooling of optically thin gas using the cooling function from Sutherland & Dopita (1993); all other parameters are as in the adiabatic simulation with a thick dense shell, and at $t=220$ and an inclination of $i=50^{\circ}$ (compare to the right panel of Figure 9). The two bipolar rings are very thin and hardly recognizable.
@@
We expect an adiabatic flow, where the radiative cooling timescale is much longer than the flow timescale, in two limits: either an optically thick flow where the photon diffusion time is much longer than the flow time, or in a low-density, optically thin gas where the radiative cooling time is much longer than the flow time. We discussed these adiabatic limits for jet-shell interaction in Akashi & Soker (2013). Scaling equations from that study, we find the condition for long radiative cooling in an optically thin flow to be that the jet-shell interaction occurs at a large distance of
@@
where $4\pi\delta$ is the solid angle covered by the two jets. We consider the condition for long radiative cooling in an optically thin flow unlikely in a rapid, violent binary interaction. Instead, we consider the jet-shell interaction to occur under conditions where the photon diffusion time is longer than the expansion time. We take $v_{f}$ as the flow velocity of the jet-shell interaction structure after the jet-shell interaction, to be somewhere below the jet’s initial velocity. The condition for adiabatic flow is that the interaction takes place very close to the AGB star, at a radius
@@
We conclude that the formation of prominent rings requires the jet-shell interaction to take place close to the binary system. In that case, the orbital motion of the binary system has a significant impact that future simulations must take into account. Namely, the interaction might take place on a time scale shorter, or not much longer, than the orbital period. Such an interaction is consistent with a bright transient event, specifically an intermediate luminosity optical transient (ILOT) event. Soker & Kashi (2012) discussed the formation of some bipolar PNe in ILOT events. The typical interaction timescale will be 0.010.001$\approx$ 0.01-0.001≈ 0.01 - 0.001 times the length of our simulations here, namely, about several months to a few years; we can scale our results concerning the shaping of rings to smaller distances and shorter times.
@@ 4.1 Some earlier studies
Many PNe possess equatorial and bipolar ring structures. Understanding the physical processes that shape these structures remains a central question in the late stages of stellar evolution. We take the approach that jets launched by a binary, or tertiary, system shape bipolar rings. There are other models. Some early theoretical frameworks demonstrated that magnetized winds and stellar rotation might produce bipolar and equatorial density enhancements. García-Segura et al. (1999) used 2D magneto-hydrodynamic (MHD) simulations to show how rotation and magnetic fields in post-AGB winds yield axisymmetric, hourglass-shaped nebulae with collimated polar outflows—structures akin to equatorial rings and bipolar lobes. Extending this to 3D, García-Segura & López (2000) explored point-symmetric features and misaligned magnetic axes. Subsequently, García-Segura et al. (2001) modeled magnetic-pressure oscillations during late AGB phases, showing how periodic field reversals generate concentric shell-like rings. Binary-oriented MHD models, such as García-Segura et al. (2018), further emphasize the interplay of jets, magnetic fields, and companion-driven dynamics.
@@ 4.2 Summary of main results
In the case of adiabatic flow interaction, we can find an inclination angle for which the two rings appear similar to the two IR rings of NGC 1514 (Figure 12). At a larger angle, namely, almost an edge-on line of sight, the two rings do not cross each other and can resemble the two rings of PN Abell 14 (lower panel of Figure 11). When radiative cooling is significant, the two rings are much less prominent and might not be observed (Figures 10 and 13). We expect the shaping to occur at an early phase, as the jet-shell forms close to the binary system and the photon diffusion time is much longer than the flow time (Section 3.2.4).
@@ Acknowledgements
A grant from the Pazy Foundation supported this research.

### v2

@@ 3.2.4 Effects of Radiative Cooling
Figure 10 of the column density maps for $\gamma=1.05$ presents at late times a bipolar ring structure, but not as prominent as in the non-radiative $\gamma=5/3$ simulations that we present in Figures 9 and 11. To better assess the impact of radiative cooling on the ring formation process, we performed an additional simulation including optically thin radiative cooling using the tabulated cooling function of Sutherland & Dopita (1993b); we kept all other initial conditions and jet parameters as in the non-radiative simulation with the thick dense shell. In Figure 13 we present the column density of this simulation at $t=220{~\rm yr}$ and an inclination angle of $i=50^{\circ}$.
Figure 13: A column-density map of a simulation with a radiative cooling of optically thin gas using the cooling function from Sutherland & Dopita (1993b); all other parameters are as in the non-radiative simulation with a thick dense shell, and at $t=220$ and an inclination of $i=50^{\circ}$ (compare to the right panel of Figure 9). The two bipolar rings are very thin and hardly recognizable.
@@
We expect a non-radiative flow, where the radiative cooling timescale is much longer than the flow timescale, in two limits: either an optically thick flow where the photon diffusion time is much longer than the flow time, or in a low-density, optically thin gas where the radiative cooling time is much longer than the flow time. We discussed these non-radiative limits for jet-shell interaction in Akashi & Soker (2013). Scaling equations from that study, we find the condition for long radiative cooling in an optically thin flow to be that the jet-shell interaction occurs at a large distance of
@@
where $4\pi\delta$ is the solid angle covered by the two jets. We consider the condition for long radiative cooling in an optically thin flow unlikely in a rapid, violent binary interaction. Instead, we consider the jet-shell interaction to occur under conditions where the photon diffusion time is longer than the expansion time. We take $v_{f}$ as the flow velocity of the jet-shell interaction structure after the jet-shell interaction, to be somewhere below the jet’s initial velocity. The condition for a non-radiative flow is that the interaction takes place very close to the AGB star, at a radius
@@
We conclude that the formation of prominent rings requires the jet-shell interaction to take place close to the binary system. In that case, the orbital motion of the binary system has a significant impact that future simulations must take into account. Namely, the interaction might take place on a time scale shorter, or not much longer, than the orbital period. Such an interaction is consistent with a bright transient event, specifically an intermediate luminosity optical transient (ILOT) event. Soker & Kashi (2012) discussed the formation of some bipolar PNe in ILOT events. The typical interaction timescale will be $\approx 0.01-0.001$ times the length of our simulations here, namely, about several months to a few years; we can scale our results concerning the shaping of rings to smaller distances and shorter times.
@@ 4.1 Some earlier studies
Many PNe possess equatorial and bipolar ring structures. Understanding the physical processes that shape these structures remains a central question in the late stages of stellar evolution. We take the approach that jets launched by a binary, or tertiary, system shape bipolar rings. There are other models. Some early theoretical frameworks demonstrated that magnetized winds and stellar rotation might produce bipolar and equatorial density enhancements. García-Segura et al. (1999) used 2D magneto-hydrodynamic (MHD) simulations to show how rotation and magnetic fields in post-AGB winds yield axisymmetric, hourglass-shaped nebulae with collimated polar outflows—structures akin to equatorial rings and bipolar lobes. Extending this to 3D, García-Segura & López (2000) explored point‑symmetric features and misaligned magnetic axes. Subsequently, García-Segura et al. (2001) modeled magnetic-pressure oscillations during late AGB phases, showing how periodic field reversals generate concentric shell-like rings. Binary-oriented MHD models, such as García-Segura et al. (2018), further emphasize the interplay of jets, magnetic fields, and companion-driven dynamics.
@@ 4.2 Summary of main results
In the case of non-radiative flow interaction, we can find an inclination angle for which the two rings appear similar to the two IR rings of NGC 1514 (Figure 12). At a larger angle, namely, almost an edge-on line of sight, the two rings do not cross each other and can resemble the two rings of PN Abell 14 (lower panel of Figure 11). When radiative cooling is significant, the two rings are much less prominent and might not be observed (Figures 10 and 13). We expect the shaping to occur at an early phase, as the jet-shell forms close to the binary system and the photon diffusion time is much longer than the flow time (Section 3.2.4).
@@ Acknowledgements
We thank an anonymous referee for useful comments. A grant from the Pazy Foundation supported this research. N.S. thanks the Charles Wolfson Academic Chair at the Technion for the support.
Software
We use the following software.
  • FLASH: A modular, parallel, adaptive-mesh, multi-physics simulation code developed at the University of Chicago, designed to model astrophysical phenomena including compressible, reactive flows (Fryxell et al., 2000).
  • VisIt: An open-source, interactive, parallel visualization and analysis tool developed by Lawrence Livermore National Laboratory, used here for post-processing and visualization of simulation outputs (Childs et al., 2012).
  • Python: A high-level programming language employed for supplementary data analysis and plotting, along with the NumPy and Matplotlib libraries (Harris et al., 2020; Hunter, 2007).
Data and Code Availability
The codes used in this article and the data underlying this article will be shared on reasonable request to the corresponding author.